### v1

# for Web Agents

Abstract
The proliferation of web agents necessitates advanced navigation and interaction strategies within complex web environments. Current models often struggle with efficient navigation and action execution due to limited visibility and understanding of web structures. Our proposed R2D2 framework addresses these challenges by integrating two paradigms: Remember and Reflect. The Remember paradigm utilizes a replay buffer that aids agents in reconstructing the web environment dynamically, thus enabling the formulation of a detailed “map” of previously visited pages. This helps in reducing navigational errors and optimizing the decision-making process during web interactions. Conversely, the Reflect paradigm allows agents to learn from past mistakes by providing a mechanism for error analysis and strategy refinement, enhancing overall task performance. We evaluate R2D2 using the WebArenabenchmark, demonstrating substantial improvements over existing methods, including a 50% reduction in navigation errors and a threefold increase in task completion rates. Our findings suggest that a combination of memory-enhanced navigation and reflective learning promisingly advances the capabilities of web agents, potentially benefiting various applications such as automated customer service and personal digital assistants.
R2D2: Remembering, Reflecting and Dynamic Decision Making
for Web Agents
Tenghao Huang1thanks: Work done during Tenghao’s internship at IBM research.  Kinjal Basu2  Ibrahim Abdelaziz2 Pavan Kapanipathi2  Jonathan May1  Muhao Chen3 1University of Southern California, 2IBM Research, 3University of California, Daivs tenghaoh@usc.edu
1 Introduction
Web agents—–autonomous AI agents designed to navigate and perform natural language-described tasks within web environments—-have become increasingly integral to applications such as online customer service Huang et al. (2024a), automated data retrieval Huang et al. (2024b), and personalized digital assistants. [^1] https://www.anthropic.com/news/3-5-models-and-computer-use. These agents interact with complex web interfaces to execute user-described tasks, often emulating human actions like clicking buttons, filling forms, and extracting information Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2023).
Figure 1: Traditional methodologies conceptualize web navigation within the framework of an Unknown MDP. The ReACT agent operates under high uncertainty due to incomplete information regarding the outcomes of their actions, leading to erroneous navigational paths, impeding effective task resolution. R2D2 transforms the task into a Known MDP, improving robustness.
Figure 2: This diagram represents various approaches for web agents framed as a search problem, where each node symbolizes a webpage. (a) Reactive: The agent chooses the best immediate actions without any proactive strategy. (b) Tree-search with reflections: the agent investigates different routes by actively navigating websites and allows for reversing direction (shown by dashed arrows). Both a) and b) approaches are Unknown-MDP-based. At each timestep, agents’ observation space is constrained, which typically results in suboptimal or inefficient results. c) R2D2: Our proposed framework constructs the search space leveraging stored state information from a replay buffer. By transforming the task into a Known MDP, R2D2 enhance agents’ ability to navigate and interact with web interfaces.
Despite recent advancements in web agents’ capabilities, a persistent challenge remains: agents frequently fail to navigate effectively within intricate web environments. As illustrated in Figure 1, the successful resolution of a user query usually requires a long series of actions.
The fundamental challenges associated with previous methodologies are twofold: first, these approaches model web navigation as an Unknown Markov Decision Process (MDP), wherein the agent has limited visibility into the consequences of its actions, often leading to suboptimal performance outcomes. Second, prior methods engage in complex reasoning during the inference phase while observing a stream of experiences, and in their simplest forms, they discard incoming trajectories immediately after a single episode. This rapid forgetting of potentially valuable experiences impedes the agent’s capacity to leverage useful information for future decision-making.
Meanwhile, a primary obstacle in enhancing web agent performance lies in navigation-related failures, which account for approximately 60% of their operational errors (as illustrated in Figure 6). These failures occur when agents become disoriented within the web environment, preventing them from reaching the target webpages necessary to execute desired tasks. Such navigation inefficiencies significantly hinder the overall effectiveness of web agents. The remaining 40% of errors stem from execution failures and edge cases, where agents either misinterpret user intentions or mishandle specific web elements.
Inspired by cognitive studies showing that humans excel at complex tasks by iteratively refining strategies based on feedback Palenciano et al. (2021); Zenkri et al. (2024), as well as by approaches in robotics for structured exploration of unfamiliar spaces Thrun (2002), we propose the R2D2 (Remembering, Reflecting, and Dynamic Decision Making) framework that enhances both navigation and task execution for web agents. Our method transforms the task from Unknown-MDP into a Known MDP by introducing the Remember Paradigm. It leverages a structured replay buffer of the agent’s experiences that guides the agent to more promising avenues Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2017); Savinov et al. (2018). At a high level, our approach enables the agent to record and recall previously visited pages—essentially constructing a dynamic map of the web environment—and then leverage this knowledge to refine its strategies. By converting the agent’s experience into a well-organized search space, we empower it to identify reliable navigation routes to target resources rather than re-deriving them from scratch during inference. This reduces the computational overhead at test time and helps avoid repetitive or unproductive exploration.
To enable continual improvement based on both successes and failures, R2D2 incorporates the Reflect Paradigm. Previous efforts in this domain often focus narrowly on immediate, execution-level errors and struggle with more pervasive navigation challenges Shinn et al. (2023); Pan et al. (2024); Wang et al. (2024a). In contrast, our method leverages the refined, known search space described earlier to minimize navigational missteps, allowing the reflection mechanism to operate more effectively on remaining execution problems. By reducing the burden of basic wayfinding, the agent’s reflective capabilities can more efficiently identify and correct subtle issues, ultimately leading to a higher overall success rate on complex web tasks.
Our proposed method diverges from traditional techniques by providing a more comprehensive and structured representation of the agent’s historical experiences. Instead of simply recalling past states Agashe et al. (2024) or relying on on-the-fly reasoning Koh et al. (2024); Zhou et al. (2024), our approach organizes the agent’s accumulated experiences into a coherent and reusable resource that effectively guides future decisions. We evaluate our approach using the WebArena benchmark Zhou et al. (2023), where it achieves substantial gains compared to baseline models, including approximately a 50% reduction in navigation errors and a threefold increase in overall task completion rates. Moreover, R2D2 outperforms state-of-the-art methods by 17%, thereby demonstrating a more robust and informed capability for executing complex web-based tasks.
2 Related Works
Enhancing Web Agents. Existing research has illuminated that language models, without intervention, struggle to express linguistic intent in formal instruction that can control an extra-linguistic environment, such as web site navigation Shi et al. (2017); Liu et al. (2018); Yao et al. (2022); Zhou et al. (2023); Deng et al. (2023). This inadequacy stems primarily from the intrinsic challenges associated with perception, strategic planning, and task execution in the intricate web environments.
To mitigate these challenges, enhancements to web agents have been categorized into two following principal strategies. (1) Perception Alignment: This strategy aims to augment agents’ capabilities in interpreting graphical user interface elements by integrating multimodal data from webpages, enhancing both textual and visual comprehension Gou et al. (2024); Anonymous (2024). (2) Post-hoc Reflection: Studies indicate that enabling agents to engage in reflective practices post interaction can facilitate learning from historical trajectories, thereby improving future task executions Shinn et al. (2023); Song et al. (2024); Pan et al. (2024); Wang et al. (2024a). (3) Online Search Algorithms: This involves the adoption of sophisticated search algorithms, including Monte Carlo Tree Search and other tree-based exploration methods, integrated with high-level planning driven by LLM-derived knowledge Zhang et al. (2024); Koh et al. (2024); Meng et al. (2024). Furthermore, Gu et al. 2024 discusses speculative planning that leverages simulations of world models.
Despite these enhancements, the performance of current web agents are constrained by the assumptions of Unknown MDP, where the potential outcomes of actions are not available. In contrast, this paper proposes a novel approach where we reconstruct the web environment’s structure based on the agents’ exploratory actions, thereby furnishing them with outcome information crucial for making informed and grounded decisions.
Continual Exploration of the Agentic Environment. Tasking agents to explore an unknown environment has been an active research direction in the community Brohan et al. (2023). Recent studies have focused on how agents abstract experiences into actionable skills, a development that is becoming increasingly central to advancements in this field Wang et al. (2023a, b); Liu et al. (2024). Within the domain of web-based agents, these skills are often conceptualized as workflows. Sodhi et al. (2024) have introduced a novel framework that leverages human-engineered workflows to compose policies to tackle web tasks. Although improving agents’ performance, manually crafting workflows can be a tedious process.
Unlike previous strategies that rely solely on high-quality successful trajectories or hand-crafted workflows, R2D2 introduces a two-part mechanism that continuously learns from the full range of agent experiences, including failed attempts. R2D2 moves beyond the limitations of purely Unknown-MDP-based assumptions and handcrafted workflows, resulting in more informed, robust decision-making and improved overall performance.
Figure 3: An overview of R2D2 architecture, highlighting the Remember and Reflect Paradigms. The Remember paradigm constructs a structured replay buffer from previous observations, enabling the agent to utilize past episodic data through A* search for navigation. Meanwhile, the Reflect paradigm diagnoses errors and generates corrective insights, which are then stored in a reflective memory for future decision-making processes.
3 Method
In this section, we present R2D2, a framework for tackling complex web navigation tasks by integrating two paradigms: Remember and Reflect. The Remember paradigm constructs a structured replay buffer from past observations(Section 3.2), while the Reflect paradigm diagnoses and corrects errors in failed trajectories(Section 3.3). We then introduce mechanisms of the reflective memory (Section 3.4). Finally, we illustrate how these paradigms interact to improve the agent’s performance through retrieval and in-context learning demonstrations (Section 3.5).
3.1 Method Overview
Given a user’s task query $q$ and an initial observation $o_{0}$ from the environment, the agent must produce a sequence of actions to address $q$. We define an episode as the process where the agent starts from $o_{0}$ and executes a trajectory $t$. Let $t=\{a_{1},a_{2},\ldots,a_{H}\}$ be the trajectory of length $H$, where each $a_{h}$ is an action at step $h$. After each action $a_{h}$, the agent receives an observation $o_{h}$, thereby forming an observation sequence $O=\{o_{1},o_{2},\ldots,o_{H}\}$. Consider $N$ distinct user queries $\{q_{1},q_{2},\ldots,q_{N}\}$, each associated with its own episode and and observation sequence $O_{i}=\{o_{i,1},o_{i,2},\ldots,o_{i,H}\}$. Across all $N$ user queries, R2D2 forms the union of these observations:$\mathcal{O}_{\text{all}}=\bigcup_{i=1}^{N}O_{i}.$
R2D2 addresses errors in trajectories by categorizing them into two distinct types: (1) Navigation failure. We define $O^{*}$ as the key observations essential for successfully addressing the query $q$. After performing the trajectory $t$, the observation sequence do not contain all the key observations, which leads to agents’ navigation failure in the environment, formally $O\cap O^{*}\neq O^{*}$. The agent fails because of incomplete information or tools to address the user query. (2) Execution failure. After performing $t$, $O\cap O^{*}=O^{*}$. In other words, navigation was successful, since the proper path was followed, but the agent still failed to address the user query.
The Remember paradigm aims to build a replay buffer from $\mathcal{O}_{all}$, allowing the agent to store and revisit past observations and experiences Blundell et al. (2016); Schaul et al. (2016); Parisotto and Salakhutdinov (2017); Savinov et al. (2018). The Reflection paradigm then corrects trajectories that failed due to execution issues, providing explicit rationales for these execution failures. The successful and corrected trajectories and their rationales are stored in a reflective memory for agent’s future reference. Finally, the Retriever leverages this reflective memory by selecting relevant corrected trajectories as in-context demonstrations, thereby continually improving the agent’s performance.
3.2 Remember Paradigm
Building the Replay Buffer. To effectively represent the web environment and enable its systematic reconstruction from the observation sequence $\mathcal{O}_{\text{all}}$, R2D2 structures the replay buffer as a directed graph $G=(O,E)$. Here, the vertex set $O$ includes the root node $o_{0}$, which corresponds to the homepage observation, and each subsequent vertex represents another webpage observation $o_{i}$. The edge set $E$ consists of tuples $((o_{i},o_{j}),a)$ where each edge corresponds to an action $a$ that transitions the agent from one observation $o_{i}$ to another observation $o_{j}$. Due to the noisy and dynamic nature of web pages, R2D2 store the differences between consecutive observations at each vertex rather than the full webpage state.
$\mathbf{A}^{*}$ Search within the Buffer. The search algorithm employs a breadth-first search ($A^{*}$) strategy Hart et al. (1968); Meng et al. (2024) to navigate and evaluate web environments effectively. Instead of expanding nodes level-by-level, R2D2 incorporates a heuristic that guides the search toward potentially relevant and promising webpages more efficiently Bonet and Geffner (1999); Guez et al. (2018); Moldovan and Abbeel (2012). This heuristic, provided by the LLM, estimates the relevance and utility of exploring a particular webpage node, thereby reducing unnecessary expansions and focusing on paths that are more likely to yield correct information.
In $A^{*}$ search, each node $o$ in the replay buffer graph is associated with both a cost (e.g., the depth from the start node or the number of steps taken) and a heuristic $h(o)$, which estimates how close $o$ is to a relevant webpage that can answer the query $q$. We derive the heuristic by prompting the LLM to assess the likelihood that the subtree rooted at $o$ will yield information relevant to $q$. $A^{*}$ search proceeds by maintaining a priority queue that selects which node to expand next based on the sum of the cost-to-come and the heuristic estimate. Webpages that are deemed relevant are added to a candidate queue, prioritizing content that potentially answers the query, while non-relevant pages are bypassed to streamline the search. This exploration continues until all reachable nodes have been evaluated. Subsequently, for each candidate node in the queue, paths are constructed back to the root, mapping feasible routes that could satisfy the query. The LLM then ranks these paths based on relevance and utility, culminating in the selection of the optimal trajectory $P^{*}$.
Algorithm 1 Optimized Web Search Using $A^{*}$ and Language Model
0:  User query $q$, Initial observation $o_{0}$
0:  Optimal solution trajectory $P^{*}$ for query $q$
1:  Initialize replay buffer graph $G=(O,E)$
2:  Add root node $o_{0}$ to $O$ as $o_{0}$
3:  Initialize priority queue $Q_{A^{*}}$ with $(o_{0},f(o_{0})=h(o_{0}))$
4:  Initialize candidate queue $Q_{cand}\leftarrow\emptyset$
5:  while $Q_{A^{*}}$ is not empty do
6:     $(o_{i},f(o_{i}))\leftarrow\text{dequeue}(Q)$
7:     if $\text{IsRelevant}(o_{i},q,LLM)$ then
8:        $Q_{cand}\leftarrow\text{enqueue}(Q_{cand},(o_{i},f(o_{i})))$
9:     end if
10:     for each action $a$ available at $o_{i}$ do
11:        $o_{j}\leftarrow\text{Transition}(o_{i},a)$Obtain next observation via action $a$}
12:        if $o_{j}$ not in Visited then
13:           $h(o_{j})\leftarrow\text{Heuristic}(o_{j},q)$
14:           $f(o_{j})\leftarrow f(o_{i})+h(o_{j})$
15:           $Q_{A^{*}}\leftarrow\text{enqueue}(Q,(o_{j},f(o_{j})))$
16:        end if
17:     end for
18:  end while
19:  Initialize trajectory set $\mathcal{T}\leftarrow\emptyset$
20:  for each $o_{i}$ in $Q_{cand}$ do
21:     $t\leftarrow\text{Backtrack}(o_{i})$Generate trajectory from $o_{0}$ to $o_{i}$ by following parent pointers}
22:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
23:  end for
24:  $P^{*}\leftarrow\text{RankAndSelectOptimal}(\mathcal{T},q)$Use LLM to rank trajectories based on relevance and utility}
25:  return  $P^{*}$
This $A^{*}$-base approach enables more informed and targeted exploration of the replay buffer. By guiding the agent through the environment with a heuristic informed by the LLM, R2D2 narrows down the search space and accelerates the discovery of relevant information. This ultimately results in faster and more accurate trajectory generation, effectively addressing complex user queries.
3.3 Reflect Paradigm
R2D2 first determines [^2] Please refer to Figure 7 in Appendices for the detailed prompt. the error type of a failed trajectory $t$. For all navigation failures, R2D2 performs search within the replay buffer to correct their navigation behaviors. We discuss details of how we address navigation failure in Section 3.2. We now discuss using reflection techniques to address execution errors.
The reflection process is designed to enhance the system’s capability to learn from mistakes within trajectories rather than only successes Madaan et al. (2023); Shinn et al. (2023). When the failure reason of a trajectory $t$ is classified as an execution failure, we prompt the LLM to identify the first erroneous action $a_{i}$. The trajectory is then truncated to include only the actions before the error point, $\{a_{1},a_{2},\ldots,a_{i-1}\}$, which are considered correct. Following this, a detailed reflection on the erroneous action $a_{i}$ is generated, providing a rationale for its failure and potential strategies for avoidance in the future. This reflection, along with the truncated trajectory, is stored in the reflective memory Weston et al. (2015); Mirowski et al. (2017); Wayne et al. (2018) that is to be introduced in Section 3.4.
3.4 Reflective Memory Mechanism
We hereby introduce the reflective memory mechanism that stores corrected and truncated trajectories for future retrieval. The reflective memory is structured as a key-value store:
Key-Value Architecture. The reflective memory mechanism functions as a key-value store where each user query is encoded into a unique query vector serving as the key, encapsulating the query’s semantic intent for efficient retrieval via vector similarity metrics. The corresponding value comprises a truncated and corrected trajectory, as described in Section 3.1, along with reflective insights. Specifically, for execution failures, only steps up to the first error are stored, while for navigation failures, corrected trajectory segments are retained once identified. During inference, a new query vector is generated and matched against existing keys to retrieve the most relevant trajectories.
Basic Operations. In alignment with conventional memory module architectures, the reflective memory mechanism defines two basic operations: (1) Lookup. Given a query, the memory retrieves the value(s) associated with the closest key vectors. (2) Update. If a newly truncated trajectory provides a more accurate or enriched reflection for an existing query, the memory updates the current value. R2D2 uses an LLM to make such decision. Please refer to Figure 8 in the Appendix for an example.
3.5 Paradigm Coordination and Inference
Exploration Phase. Using a ReACT agent Yao et al. (2023), R2D2 processes user queries and collects observational data to build $\mathcal{O}_{\text{all}}$. Trajectories are classified and corrected via Remember and Reflect paradigms, then stored in the memory.
Inference Phase. During inference, user queries are encoded into vectors and matched against reflective memory to retrieve relevant trajectories as in-context demonstrations Karpukhin et al. (2020); Brown et al. (2020). These demonstrations guide the agent’s response. Failed trajectories undergo reflection, and the memory is updated to improve future performance. This coordination allows R2D2 to leverage past experiences and reflections, ensuring continuous learning and enhanced handling of complex queries.
4 Experiments
In this section, we evaluate the proposed R2D2 framework for web agent tasks and compare it with baseline methods. We first delve into the details of our experimental setup (section 4.1), discuss the results obtained(section 4.2), and perform ablation studies to understand the strengths of different components (section 4.3). Furthermore, we provide a comprehensive error analysis (Section 4.4).
4.1 Experimental Setup
Method Tasks Total SR
CMS Reddit Shopping Map GitLab
ReACTYao et al. (2023) 13.1%
Tree-Search Koh et al. (2024) 16.5% 10.5% 28.1% 25.8% 13.3% 19.2%
AutoEval Pan et al. (2024) 20.2%
LATS Zhou et al. (2024) 15% 25% 30% 27% 17% 22.5%
AR Wang et al. (2024a) 16% 24% 32% 27% 18% 23.4%
BrowserGym Drouin et al. (2024) 23.5%
R2D2 30.4% 20.8% 35.8% 29.6% 28.0% 27.5%
Table 1: Performance comparison across multiple web-based tasks. Reported success rates (SR) are organized by model and method, including baseline approaches and our proposed R2D2. Superscripts indicate the model used: GPT-4o, GPT-4. The baseline results are from corresponding papers.
Benchmark. We use the WebArena benchmark Zhou et al. (2023). This benchmark comprises diverse web interaction scenarios, ranging from web shopping to customer relationship management system (CMS). The dataset consists of 812 user queries with annotated ground truth trajectories. The Webarena benchmark further provides a set of validators to programmatically validate the functional correctness of each task.
Implementation Details. We choose gpt-4o [^3] OpenAI. (2024). ChatGPT (November 20th version). as our main LLM for both Remember and Reflect paradigm. We use the retriv [^4] https://github.com/AmenRa/retriv framework as the backbone of the reflective memory index, and select “sentence-transformers/all-MiniLM-L6-v2” as the dense embedding model for our retriever.
Baselines. We compare our framework against several represetative agent frameworks: (1) ReACT Yao et al. (2023): a widely-used framework, which takes an observation of the environment as input, performs Chain-of-Thought reasoning, and then generates the next action. (2) Tree-Search Koh et al. (2024): an inference-time tree-search strategy to perform best-first tree-search in web environments. It enables agents to revert to the most recently validated state upon encountering a failed trajectory. (3) LATS Zhou et al. (2024): a method based on Monte Carlo tree search that employs LLMs as agents, value functions, and optimizers for decision-making. (4) Anticipatory Reflection Wang et al. (2024a): a method that explicitly considers potential failures before action, alignment and backtracking after actions to maintain task objectives. (5) AutoEval Pan et al. (2024): methods that boost agent performance using domain-general automatic evaluators. (6) BrowserGym Chezelles et al. (2024): a framework that incorporates additional actions and observation tools for agents to interact with the environment. [^5] We also notice there are other methods that use different setups, such as SteP Sodhi et al. (2024) that employs human-engineered workflows, and AWM Wang et al. (2024b) that uses BrowserGym framework that customizes a larger action space than standard WebArena to interact. For the sake of direct comparison, these frameworks that depend on additional human efforts are not taken into comparison here.
4.2 Main Results
Outperforming Tree-search and ReACT Methods:
As illustrated in Table 1, the R2D2 model achieves superior success rates across all tasks when compared to both the Tree-search and ReACT methods, demonstrating its effectiveness. For instance, in the CMS and Reddit tasks, R2D2 outperforms Tree-search by substantial margins (30.4% vs. 16.5% in CMS and 20.8% vs. 10.5% in Reddit). The improvement suggests that the integration of a systematic replay buffer and a reflective memory paradigm in R2D2 enhances the agent’s ability to navigate and learn from past interactions, avoiding repeated errors and optimizing the decision-making process based on agents’ experiences. This highlights the advantage of R2D2’s enhanced observability in the MDP, which enables a more comprehensive understanding of the web environment, leading to more accurate and efficient strategies.
Substantial Improvements in CMS and GitLab Domains:
The results in the CMS and GitLab domains are particularly notable. R2D2 achieved a 30.4% success rate in CMS and 28.0% in GitLab, considerably higher than other tested methods. These domains often require complex navigations with web interfaces, where R2D2’s capability to leverage past visited states and reflect on past actions proves especially beneficial.
Comparison with Complex Frameworks Like AR, LATS and AutoEval:
When compared to complex frameworks employing sophisticated reflection mechanisms (e.g., AR, AutoEval), R2D2 holds its own or exceeds performance, with a total success rate (SR) of 27.5%. While AR and AutoEval offer robust reflection capabilities, R2D2’s integrated approach to first remembering and then reflecting allows it to preemptively correct paths and further refine strategies. The success can be attributed to the method’s dual-paradigm system. We show more analysis in Section 4.3 and Section 5.
4.3 Ablation Study
Ablating rounds of execution. To better understand the strength of our proposed framework, we compare R2D2 with advanced baselines that emphasize reflection techniques. Figure 4 illustrates a marked increase in the success rate of R2D2 during initial episodes. Upon manual inspection, we attribute this early performance enhancement primarily to the effective resolution of navigation failures.
By the fifth episode, R2D2 substantially outperforms AR and LATS, confirming its methodological superiority. This highlights R2D2’s ability to leverage historical data and adaptive strategies effectively. While AR demonstrates commendable learning capabilities through its anticipatory reflection, it fails to match R2D2’s effectiveness. LATS, in contrast, shows minimal improvement. These findings support the practical superiority of the R2D2 model in dynamic learning environments.
Figure 4: Performance comparison with different reflection-based methods. R2D2 achieved marked increase at the first episode. Our manual inspection in Figure 6 shows 75% of the initial increase is attributed to fixing the navigation failures.
Figure 5: Performance comparison with ablation variants. Removing navigation or reflection capabilities from R2D2 is very harmful to performance.
Ablating Remember & Reflect paradigms. In this ablation study focused on the CMS domain, the full R2D2 model substantially outperforms its variants, as shown in Figure 5. The “– Reflection” variant, which lacks advanced reflection capabilities, shows moderate gains, while the “– Navigation” variant, which removes navigation, achieves only marginal improvement. Notably, the “– Reflection” variant, though initially showing some improvement, demonstrates a limited performance increase in later episodes, suggesting that while navigation capabilities can provide early benefits, their effectiveness without reflection support plateaus quickly. This observation highlights the critical role of navigation in sustaining performance improvements over time, reinforcing that reflection alone is insufficient for long-term success in complex web environments.
Ablating Failed Trajectories. To elucidate the learning dynamics of R2D2, we conduct a study to isolate the impact of failed trajectories. During this study, only successful trajectories are provided as in-context demonstrations at inference time, thereby restricting R2D2 to learning exclusively from positive examples. This variant falls 7.5% from the full implementation of R2D2 to 20.5%. This also reveals a critical limitation: the number of positive examples is insufficient to provide robust navigation and reflection to the agent during inference. Consequently, if no relevant successful trajectory is identified at retrieval time. These findings substantiate the hypothesis that failed trajectories, despite not directly addressing user queries, are instrumental in enriching R2D2’s strategic repertoire, and R2D2 extend beyond the mere memorization of positive examples.
4.4 Error Analysis
As shown in Figure 6, we manually inspect the trajectories of the same 60 queries executed by the vanilla ReACT agent and R2D2 agent. About 60% of the vanilla ReACT agent trajectories stall at the navigation stage, so there is not even an opportunity to fail in execution. In contrast, the R2D2 agent substantially reduces navigation failures, reliably guiding itself toward the right content and thereby reaching a point where it is possible to fail in execution more frequently. As a result, R2D2 achieves a higher pass rate overall. We further annotate and discuss erroneous trajectories in Appendix A.
Figure 6: Error analysis of vanilla ReACT agents’ and R2D2’s trajectories. “E” indicates execution failures, “N” indicates navigation failures, and “S” indicates success. R2D2 substantially reduces navigation failures, achieving higher success rate.
5 Qualitative Study
We present a case study to conduct a qualitative evaluation of the R2D2 framework, as detailed in Appendix C. Our framework uses precise navigation and advanced reflective capabilities, enabling effective interaction with and extraction of information from intricate web environments. In contrast, the baseline struggles with navigation and data analysis, resulting in incomplete and inaccurate outcomes. These findings emphasize the importance of meticulous navigation strategies and reflective processing for successful information retrieval in sophisticated web interfaces.
6 Conclusion
The R2D2 framework significantly enhances web agents’ capabilities by integrating Remember and Reflect paradigms, enabling more effective navigation and interaction in complex web environments. This approach leads to measurable improvements in performance, reducing errors and increasing task completion rates. Through structured memory use and reflective learning, R2D2 not only outperforms existing models but also offers a scalable solution adaptable to various domains. Future work could extend its application, further optimizing agent functionality across broader scenarios.
Limitations
Our experiments were exclusively conducted in English. This limitation restricts our understanding of the model’s efficacy across different linguistic contexts, potentially overlooking cultural and language-specific nuances that could affect the agent’s performance in non-English web environments.
Appendix A Error Analysis
Among all the execution failures of R2D2, the errors can be classified as following:
Pessimistic Reflection (30.3%). When the agent makes a mistake and enters the reflection phase, it occasionally produces overly pessimistic rationales. Instead of proposing a plausible alternative action or a corrective step—such as trying a different button or re-verifying information on the same page—the agent may hastily conclude that the service is unavailable, broken, or that no solution exists. This pessimism not only mischaracterizes the underlying issue but also inhibits effective learning from the mistake. By prematurely giving up, the agent misses opportunities to refine its approach, explore subtle variations in the action sequence, or simply retry a failed step with slight modifications.
Lack of GUI understanding (24.2%). In certain scenarios, the agent struggles to properly interpret or interact with the graphical user interface (GUI) elements of the webpage. For example, when the user’s query requires submitting information through an online form, the agent may fail to pinpoint the correct input fields or submission buttons, even after correctly navigating to the right page. As a result, it may click on the wrong element, repeatedly fail to submit required information, or get stuck trying to identify how to move forward.
Difficulty with Executing Complex Plan (20.2%). After reaching the desired section of a website, the agent may still falter when asked to carry out intricate, multi-step tasks. For instance, it could be instructed to iterate through a list of items, adding each one to a cart, verifying their details, and then proceeding to a checkout process. This difficulty suggests that, although navigation is now more reliable, the agent still needs improved reasoning capabilities and better long-term action planning to handle scenarios that demand careful step-by-step execution.
Appendix B Prompt Details
In this section, Figure 7 and Figure 8 show prompt details in the paradigms.
User Query gpt-4o Failed Traj. R2D2 Intermediates R2D2 Final Traj.
Get the billing name of the oldest complete order. 1. Click ’SALES’ icon 2. Click order_id/65 3. Scroll down 4. Stop action: Grace Nguyen
5. [Uncaptioned image]
ROUND 1
Navigation: 1. Click ’SALES’ 2. Click ’Orders’
Reflection: Agent fails to reach the correct page.
1. Click ’SALES’ icon 2. Click ’Orders’ 3. Click ’↑ Purchase Date’ 4. Click ’Filters’ icon 5. Click ’Status’ 6. Type ’Complete’ 7. Click ’Apply Filters’ 8. Click ’View’ 9. Stop action: John Lee
10. [Uncaptioned image]
List the top 2 search terms in my store. 1. Click ’CATALOG’ 2. Click ’SALES’ 3. Click ’REPORTS’ 4. Click ’CATALOG’
5. [Uncaptioned image]
ROUND 1
Navigation: 1. Click ’REPORTS’ 2. Click ’Search Terms’
Reflection: Agent fails to reach the correct page.
ROUND 3
Navigation: Correct navigation!
Reflection: Arrives correctly but needs to sort entries.
1. Click ’REPORTS’ 2. Click ’Search Terms’ 3. Click ’Hits’ 4. Click ’↓ Hits’ 5. Click ’↑ Hits’ 6. Stop action: Top 2 search terms are ’hollister’ (19 hits) and ’Joust Bag’ (10 hits).
7. [Uncaptioned image]
Table 2: Comparison of action trajectories for resolving user queries. This table illustrates how the GPT-4o model fails to navigate complex web interfaces, whereas the proposed R2D2 framework successfully manages both intermediate steps and final actions to accurately address the same queries. Please refer to Section 5 for more detailed analysis.
Figure 7: Error classification prompt to determine agent trajectory error type.
Figure 8: Trajectory evalution prompt for Update Operation. The LLM compares two trajectory and determines which one is better for addressing the user query.
Appendix C Qualitative Analysis
As shown in Table 2, an illustrative case can be observed in the user query asking for the billing name of the oldest complete order. The trajectory implemented by gpt-4o failed to resolve the query as it did not properly navigate the complex web interface. In contrast, the trajectory of R2D2 shows a series of steps that meticulously navigate through the interface. This targeted navigation led directly to the successful completion of the task, emphasizing the necessity of precise and thoughtful navigation strategies to effectively interact with and extract information from sophisticated web environments.
In instances where navigation is executed correctly but is insufficient to solve the task, the reflection module of R2D2 plays a crucial role. A clear example of this is the task to list the top 2 search terms in the store. While the gpt-4o trajectory navigates correctly to the ’Search Terms’ section, it does not delve deeper into analyzing or sorting the data, resulting in incomplete and inaccurate information retrieval. Conversely, R2D2 not only accesses the correct section but also actively manipulates the data display by sorting the search terms according to their hits, thereby precisely identifying and articulating the top search terms. This demonstrates the power of R2D2’s reflective capabilities.